OpenStudy (anonymous):

What is the product of the polynomials below? (7x2 + 9x + 8)(9x3 + 7x) A. 63x6 + 81x4 + 121x3 + 63x2 + 56x B. 63x5 + 153x4 + 112x3 + 56x C. 63x5 + 81x4 - 121x3 + 63x1 + 56x D. 63x5 + 81x4 + 121x3 + 63x2 + 56x

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (7x^2 + 9x + 8)(9x^3 + 7x) }\) \(\Large \color{MidnightBlue}{\Rightarrow 9x^3(7x^2 + 9x + 8) + 7x(7x^2 + 9x + 8) }\)

OpenStudy (anonymous):

would that be c

OpenStudy (anonymous):

hello I need HELP

OpenStudy (pfenn1):

No, it isn't C. Did you multiply it out and collect the terms?

OpenStudy (anonymous):

YEAA I DID

OpenStudy (pfenn1):

So how could you have gotten negative terms as shown in C when all your original multipliers are positive?

OpenStudy (anonymous):

WELL ION KNOW HOW TO DO IT ' MUST BE DID SUMN WRONG

OpenStudy (pfenn1):

ok, let's start with what @ParthKohli showed up above. The expression can be written as\[{ 9x^3(7x^2 + 9x + 8) + 7x(7x^2 + 9x + 8) }\]Let's just multiply out everything. what is \[9x^3 \times 7x^2\]?

OpenStudy (pfenn1):

hint \[x^a \times x^b=x^{a+b}\]